1. A frustum of a regular hexagonal pyramid has an upper base edge of 5m and a lower base edge of 8.5m. Its lateral area is 160m2. Determine the slant height of the frustum.

2. A regular square pyramid has an altitude of 12 m and volume of 196 m3. Determine the slant height of the pyramid.

3. A regular square pyramid has a height of 8m. If the slant height makes an angle of 45º with the base, find the lateral
area of the pyramid
1
Expert's answer
2020-12-13T17:42:54-0500



Its lateral are is equal to 160m2 and consists 6 trapezoid which has an upper base edge of 5m and a lower base edge of 8.5m. we can see in the picture above the slant height of the frustum is the altitude of the equilateral trapezoidal height at the side surface:



The are of this trapezoid is "\frac{160}{6}" ="\frac{80}{3}" . The formula to find are of trapazoid is "S=\frac{x+y}{2}*h" .Here x=5m, y=8.5 m. So:

h="\frac {2*S}{x+y}" ="\frac{2*\frac{80}{3}}{5+8.5}=" 3.95 (m)



2.


This picture explain our exercise. We have V=196 m3, h = 12 m, b-? .

Firstly, to find b, we find a. We use formula of finding volume: "V=\frac{1}{3}" * a2 * h.

a="\sqrt{\frac{3*V}{h}}" = "\sqrt{\frac{3*196}{12}}" = 7 (m).

Then we have only to find b. We can use pythagor theorem by this picture:



b ="\sqrt{x^2+y^2}" = "\sqrt{12^2+7^2}=13" m.



3.


This picture explain our exercise. We have h =8m.

The angle between base and slant height is 45o. So h="\frac{a}{2}" ---> a=2*h=16 m, and b = "h*\sqrt2" ="8*\sqrt2"


Lateral area consists 4 equilateral triangles like this:



We will compute one of these triangles and we will multiply by 4.

The formula to find the area of triangle which is above: "S=\frac{a \times h}{2}" ;

Here is a, h=b;

"S_1=\frac{a\times b}{2}=\frac{16\times 8\times \sqrt2}{2}=64\sqrt2;"

S"_{overal}=4\times 64\times\sqrt2=256\sqrt2;"